# Learning to Predict Structural Vibrations

### v3

@@ Appendix C Architectures
Table 5: Model size and speed comparison for a forward pass of a batch of 32 plate geometries on an A100 GPU. In comparison, solving one geometry via FEM takes 2 minutes and 19 seconds. The slowest deep learning method is then around 6000 times faster.
@@
FQO-RN18 12.7 0.006
FQO-ViT 9.28 0.010
Grid-UNet 27.9 0.038
FQO-UNet 7.1 0.669
FEM (20 CPU cores) - $\sim$ 4448.000
@@ Appendix D Ablations
We provide ablation results for the $\alpha$ value for calculating the total training loss $L_{total}=\alpha L_{\mathbf{V}}+(1-\alpha)L_{F}$ in Table 6. We find that the loss on the velocity field prediction $L_{\mathbf{V}}$ is more important than the loss on the frequency response.
@@
We provide ablation results on the number of channels in the FQO-UNet and Grid-UNet architectures in Table 7. We find that increasing the number of channels did not lead to a perfomance improvement for the Grid-UNet and leads to marginal further improvements for the FQO-UNet architecture.
@@ E.1 Sample Efficiency
10 % 0.73 19.44 0.49 7.9 0.56 21.67 0.42 10.0
25 % 0.58 14.27 0.31 7.2 0.43 18.17 0.42 7.0
50 % 0.45 12.20 0.20 6.4 0.31 14.31 0.32 5.6
75 % 0.37 10.96 0.18 5.5 0.26 10.58 0.20 5.1
Full dataset 0.32 10.70 0.17 5.3 0.22 10.37 0.16 4.8
@@ E.2 Multiple Trainings with Random Splits
To provide a notion of the variability of the results for different training runs, we performed four trainings for the FQO-UNet with different initial seeds and different random splits in training and validation set. In Table 9 we report evaluation results on the respective validation sets and on the unseen test set and observe only modest variation.
@@ E.3 Visualizations
Figure 11: G-5000 example predictions. The velocity fields at the three peaks with the highest amplitude are shown. The plots are scaled with respect to the maximum velocity in the prediction and reference velocity field to make the differences in magnitude visible.

### v4

@@ Appendix C Architectures
Table 7: Model size and speed comparison for a forward pass of a batch of 16 plate geometries on an A100 GPU. In comparison, solving one geometry via FEM takes 2 minutes and 19 seconds. The slowest deep learning method is then around 6000 times faster.
@@
FQO-RN18 12.7 0.005
FQO-ViT 9.28 0.013
Grid-UNet 27.9 0.010
FQO-UNet 7.1 0.338
FEM (20 CPU cores) - $\sim$ 2224.000
@@ C.5 Training
Compute resources.
All trainings reported in this work were computed on a cluster with single A100 GPUs. The most resource intense training run took roughly 1d and 16h on a single A100 GPU and was the ablation of the number of channels with the highest scaling factor for the FQO-UNet method detailed in Section D. All other training runs with the FQO-UNet were completed in less than 24h. The trainings for the other methods were substantially shorter with i.e. the Grid-UNet finishing training in roughly 2h - 3h and likewise the FNO (velocity field) method. The roughly 20 to 30 training runs for the FQO-UNet dominate the required compute resources. We estimate that it took in total 1 A100 for 30 days. In addition, preliminary and failed experiments required further computational resources.
@@ Appendix D Ablations
We provide ablation results for the loss function for training methods that predict the velocity field. We consider the loss function $L_{total}=\alpha L_{\mathbf{V}}+(1-\alpha)L_{F}$ and provide results in Table 8. This ablation was performed with training batches consisting of 300 frequencies per geometry, instead of a subset. We find that the loss on the velocity field prediction $L_{\mathbf{V}}$ is more important than the loss on the frequency response.
@@
We provide ablation results on the number of channels in the FQO-UNet and Grid-UNet architectures in Table 9. For the FQO-UNet, this ablation was performed with training batches consisting of 300 frequencies per geometry, instead of a subset. We find that increasing the number of channels did not lead to a perfomance improvement for the Grid-UNet and leads to marginal further improvements for the FQO-UNet architecture.
@@ E.1 Sample Efficiency
10 % 0.73 19.44 0.49 7.9 0.65 27.57 0.55 13.9
25 % 0.58 14.27 0.31 7.2 0.45 21.04 0.44 8.9
50 % 0.45 12.20 0.20 6.4 0.35 17.00 0.15 7.2
75 % 0.37 10.96 0.18 5.5 0.29 15.06 0.15 5.9
Full dataset 0.32 10.70 0.17 5.3 0.24 13.51 0.13 5.1
Full results on training with different ratios of frequencies and geometries are provided in Table 11.
Table 11: Data generation experiment.
V-5000
# Freqs # geometries $\mathcal{E}_{\text{MSE}}$ $\mathcal{E}_{\text{EMD}}$ $\mathcal{E}_{\text{PEAKS}}$ $\mathcal{E}_{\text{F}}$
300 500 0.48 13.16 0.32 5.7
150 1,000 0.31 11.18 0.22 4.3
30 5,000 0.12 8.74 0.14 2.1
10 15,000 0.10 11.18 0.17 1.6
3 50,000 0.10 11.47 0.20 1.4
15 50,000 0.02 4.06 0.04 0.08
original 300 5,000 0.08 4.24 0.07 1.7
@@ E.2 Multiple Trainings with Random Splits
To provide a notion of the variability of the results for different training runs, we performed four trainings for the FQO-UNet with different initial seeds and different random splits in training and validation set. These trainings were performed with training batches consisting of all 300 frequencies per geometry, instead of a subset. In Table 12 we report evaluation results on the respective validation sets and on the unseen test set and observe only modest variation.
@@ E.3 Visualizations
Figure 12: G-5000 example predictions. The velocity fields at the three peaks with the highest amplitude are shown. Empty axes indicates less than three peaks in the response. The plots are scaled with respect to the maximum velocity in the prediction and reference velocity field to make the differences in magnitude visible.